OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2 ?

OpenStudy (anonymous):

\[Log_2 x- Log_2 4=2\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[Log_2 4=a\] same as asking \[2^a=4\] two to what raised equals 4? I think 2^2=4 so a=2 we may replace \[Log_2 4\] with 2

OpenStudy (anonymous):

\[Log_2 x-2=2\] \[Log_2 x=4\] \[2^4=x\] x=16

OpenStudy (anonymous):

oh ok..thanks for showing the work...I actually just learned how to to do this

OpenStudy (anonymous):

thanks for being interested in work, that's rare around here

OpenStudy (anonymous):

lol